A body of mass m released from height H .the avrage power for complete motion will be

A  body of mass m released from height H .the avrage power  for complete motion will be

Grade:12th pass

1 Answers

Khimraj
5 years ago
work done = mgH
time taken = \sqrt{2H/g}
average power = W/T = mg\sqrt{gH/2}
Hope it clears...............................................................